The junior homecoming float needs a total of 600 flowers for the base. The junior
class has already made 200 flowers. Lacy and Sydney need to finish making paper
flowers. Lacy makes 6 flowers every minute and Sydney makes 4 flowers every
minute. Both girls work for the same amount of time. Write an equation , that will help find the number of minutes, m, it takes Lacy and Sydney to make the remaining
flowers?

Answer by ikleyn(52818)

.
The junior homecoming float needs a total of 600 flowers for the base. The junior
class has already made 200 flowers. Lacy and Sydney need to finish making paper
flowers. Lacy makes 6 flowers every minute and Sydney makes 4 flowers every
minute. Both girls work for the same amount of time. Write an equation , that will help
find the number of minutes, m, it takes Lacy and Sydney to make the remaining
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

6m + 4m = 600 - 200

   10m  =    400

     m  =    400/10 = 40 minutes.    ANSWER

Solved.